LOVEMEE:

find the ratio of i) 60 to 48 ii) 3.75 kg to 750 gm

MysticalDimpless:

i) Required Ratio = \[\frac{ 60 }{ 48}\] = \[\frac{ 5}{ 4 }\] = 5 : 4 ii) Since 3.75 kg = 3.75 x 1000 gm = 3750 gm Required Ratio = \[\frac{ 3.75kg }{ 750gm }\] = \[\frac{ 3750gm }{ 750gm }\] = \[\frac{ 5 }{ 1 }\] = 5 : 1

LOVEMEE:

thnx

Astrid1:

Are you going to close this ma'am?

LOVEMEE:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Astrid1 Are you going to close this ma'am? \(\color{#0cbb34}{\text{End of Quote}}\) ok

itsme:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MysticalDimpless i) Required Ratio = \[\frac{ 60 }{ 48}\] = \[\frac{ 5}{ 4 }\] = 5 : 4 ii) Since 3.75 kg = 3.75 x 1000 gm = 3750 gm Required Ratio = \[\frac{ 3.75kg }{ 750gm }\] = \[\frac{ 3750gm }{ 750gm }\] = \[\frac{ 5 }{ 1 }\] = 5 : 1 \(\color{#0cbb34}{\text{End of Quote}}\) gud explanation gud job !!!

MysticalDimpless:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @itsme \(\color{#0cbb34}{\text{Originally Posted by}}\) @MysticalDimpless i) Required Ratio = \[\frac{ 60 }{ 48}\] = \[\frac{ 5}{ 4 }\] = 5 : 4 ii) Since 3.75 kg = 3.75 x 1000 gm = 3750 gm Required Ratio = \[\frac{ 3.75kg }{ 750gm }\] = \[\frac{ 3750gm }{ 750gm }\] = \[\frac{ 5 }{ 1 }\] = 5 : 1 \(\color{#0cbb34}{\text{End of Quote}}\) gud explanation gud job !!! \(\color{#0cbb34}{\text{End of Quote}}\) thnxx

itsme:

yw